# Involutive (simple) latin solutions of the Yang-Baxter equation and related (left) quasigroups

### v1

@@ 5.3 Simple cycle sets with regular displacement group
Now, we can give the desired description.
Theorem 44.
Let $(p,G,\phi,\psi,\alpha)$ be a quintuple such that:
@@
    $(\phi,\psi)\in Sym(G)\times Aut(G,+)$ with $\psi$ having order $p^{s}$ for some $s\in\mathbb{N}$ and such that Eq. 7 holds for all $x,y\in G$;
  • 4.
    every normal subgroup $H$ of $G$ invariant under $\psi$ is not invariant under $\phi$;
  • 5.
    $\alpha\in Aut(G,\cdotp)$, where $x\cdotp y:=\phi(x)+\psi(y)$ for all $x,y\in G$.
Then, the pair $(X,\cdotp)$ given by $X:=G$ and
$x\cdotp y:=\alpha(\phi(x)+\psi(y))$
for all $x,y\in G$ is a non-trivial simple cycle set with $\mathrm{Dis}(X)\cong G$. Conversely, every simple cycle set $X$ having nilpotent regular displacement group can be constructed in this way.
@@
If $(p,G,\phi,\psi,\alpha)$ is a quintuple as in the first part, by a standard calculation one can show that $(X,\cdotp)$ is an irretractable cycle set with $\mathrm{Dis}(X)\cong G$. By 34, this cycle set is simple.
Conversely, suppose that $X$ is a simple cycle set with nilpotent regular displacement group. By 41 and 33, $\mathrm{Dis}(X)$ is a $p-$group, for a prime number $p$, and $X$ is right linear over $\mathrm{Dis}(X)$. Moreover, by 37, the $p$-subgroup of $\mathcal{G}(X)$ is normal, hence by 22 $X$ is a deformation of a cycle set $Y$ with $\mathcal{G}(Y)$ a $p$-group. According to 29 $Y$ must be right linear, so $Y$ can be constructed by $(\phi,\psi)\in Sym(\mathrm{Dis}(X))\times Aut(\mathrm{Dis}(X))$ with $\psi$ of order $p^{s}$, for some $s\in\mathbb{N}$. If $\alpha\in Aut(Y,\cdotp)$ is such that $X=Y_{\alpha}$, we have that $X$ can be constructed by the quintuple $(p,\mathrm{Dis}(X),\phi,\psi,\alpha)$, where by simplicity of $X$ and 34 it follows that every normal subgroup of $G$ invariant under $\psi$ is not invariant under $\phi$. Therefore the statement follows. ∎
@@ 6 Affine cycle sets
The following result, that is a reformulation of [8, Theorem $4.18$], allows to detect the class of affine cycle sets by group-theoretic properties.
@@
    Let $(X,\cdotp)=\mathrm{Aff}(A,\phi,\psi,c)$ be a latin affine cycle set with underlying abelian group $(A,+)$. As showed in [8, Section $4$], for affine cycle sets Eq. 5 is equivalent to the relation $[\psi,\phi^{-1}]=id_{X}$, that is also the defining relation of the first Weyl Algebra $A_{1}(\mathbb{Z})$. Thus $(X,+)$ has a structure of $A_{1}(\mathbb{Z})$-module defined by $\rho:A_{1}(\mathbb{Z})\longrightarrow\operatorname{End}{(A,+)}$ with $\rho(a)=\psi$ and $\rho(b)=\phi^{-1}$.
@@
Note that, if $(M,+)$ is an elementary abelian $p$-group, then $(M,+)$ is a $\mathbb{Z}/p\mathbb{Z}$-module and so $\mathrm{Aff}(M,\phi,\psi,c)$ is a $A_{1}(\mathbb{Z}/p\mathbb{Z})$-module. Hence we have a representation of the algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ over $M$, and so the dimension of the representation must be a multiple of $p$ [8, Proposition 4.7].
@@ 6.1 Simple affine cycle sets
Let $p$ be a prime number. As a main consequence of 3 and 46, we obtain that every finite simple affine cycle set corresponds to a finite-dimensional irreducible representations $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ together with a constant $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, where $n:=dim(\rho)$.
@@
Conversely, suppose that $X$ is an affine simple latin cycle set. Then, by [8, Section 4] and 46 there exist a natural number $n$, a prime number $p$ dividing $n$, an element $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, and two automorphism $\phi$ and $\psi$ of $(\mathbb{Z}/p\mathbb{Z})^{n}$, such that $X:=(\mathbb{Z}/p\mathbb{Z})^{n}$ and $x\cdotp y:=\phi(x)+\psi(y)+c$ for all $x,y\in(\mathbb{Z}/p\mathbb{Z})^{n}$. Then, the assignments $a\mapsto\psi$ and $b\mapsto\phi^{-1}$ gives rise to a representation $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ (see also Remark 3). By simplicity of $X$, the irreducibility of $\rho$ follows. ∎
We show that the construction of all the simple affine cycle sets can be reduced to the complete description of the affine ones having prime-power permutation group.
Theorem 48.
Let $(p,\rho,c,\eta,g)$ be a quintuple such that:
@@
    $\rho$ is an irreducible representation of the Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ with $\rho(a):=\psi$ and $\rho(b):=\phi$ invertible matrices, and $\psi$ having order $p^{k}$ for some $k\in\mathbb{N}$;
@@
    $(\eta,g)\in C_{Aut(G,+)}(\phi,\psi)\times(\mathbb{Z}/p\mathbb{Z})^{n}$ is such that $\eta(c)-c=\phi^{-1}(g)+\psi(g)-g$.
Then, the pair $((\mathbb{Z}/p\mathbb{Z})^{n},\cdotp)$ with
$x\cdotp y:=g+\eta(\phi^{-1}(x)+\psi(y)+c)$
@@
Conversely, every non-trivial simple cycle set $X$ having a displacement group abelian and normal in $\mathcal{TG}(X)$ can be constructed in this way.
Proof.
The proof is similar to the one of 44. ∎
@@ 6.2 Simple affine cycle sets of minimal size
By 47, several examples of simple cycle sets can be obtained by irreducible representations of the first Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$. Even if these representations are not classified over arbitrary fields (to our knowledge), several examples can be recovered from the irreducible representations of the first Weyl Algebras on algebrically closed fields, that are completely classified. Indeed, if $K$ is an algebrically closed field of characteristic $p>0$, all the finite dimensional irreducible representations of $A_{1}(K):=K[a,b]/(ab-ba-1)$ are given by the $p\times p$ matrices
@@
By 47 and 50, we are able to construct concretely all the affine simple cycle sets of size $p^{p}$, for an arbitrary prime number $p$. The next goal is to distinguish the isomorphism classes.
@@
According to 47 and 50, we have that $X$ is given by $X(\lambda,\mu,c)=\mathrm{Aff}(\mathbb{Z}_{p}^{p},(\lambda I+\Delta)^{-1},\mathcal{A}_{\mu},c)$. If $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic then the matrices $A_{\mu}$ and $A_{\mu^{\prime}}$ (resp. $\mu I+\Delta$ and $\mu^{\prime}I+\Delta$) are conjugate according to 53. Therefore such matrices have the same eigenvalues, i.e. $\lambda=\lambda^{\prime}$ (resp. $\mu=\mu^{\prime}$). By this fact and 53, $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic if and only if and only if $\lambda=\lambda^{\prime}$, $\mu=\mu^{\prime}$ and there exists $\alpha\in GL_{p}(\mathbb{Z}/p\mathbb{Z})$ and $u\in\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$ such that $\mu I+\Delta=\alpha(\mu I+\Delta)\alpha^{-1}$, $\mathcal{A}_{\mu}=\alpha\mathcal{A}_{\mu}\alpha^{-1}$ and $c^{\prime}=\alpha(c+u)$. By virtue of Lemma 52 then $\alpha=sI$ with $s\neq 0$. Therefore
@@ 7 Applications and examples
We start by a non-existence result that extend [8, Lemma 4.13]. In particular, we show that we have no latin cycle sets with cyclic displacement group.
@@
Let $n:=|X|$ and suppose that $\mathrm{Dis}(X)$ is a cyclic group. At first, suppose that $n$ has an odd prime divisor $p$ and let $m$ be the positive integer such that $n/m=p$. As in 42, the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$ and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $H$ induces a congruence $\sim$ on $X$ such that $|X/\sim|=p$, and this implies that there exist a latin cycle set of size $p$, but this contradicts [24, Theorem 2.13]. Now, we consider the case $n=2^{t}$ for some natural number $t$. The case $t=1$ is not possible by [24, Theorem 2.13]. The case $t=2$ is excluded by an inspection of small cycle sets. The case $t>2$ can be showed as in the first part by [2, Proposition 5.4]. ∎
By [21] every indecomposable and irretractable cycle set of size $p^{2}$ is simple. A complete description of the latin ones follows by [21, Theorem $5.1$], as we show in the following result.
@@
By 3, $X$ must be indecomposable and irretractable. If $p=2$, we know by [8] that the two indecomposable and irretractable cycle sets of order $4$ are latin. If $p>2$, by [21, Theorem $5.1$] $X$ is given by $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ and $(a,x)\cdotp(b,y)=(\alpha b+\alpha x,\alpha y+\alpha\phi(b-a))$, where $\alpha\in\mathbb{Z}/p\mathbb{Z}\setminus\{0\}$ and $\phi$ is a map from $\mathbb{Z}/p\mathbb{Z}$ to $\mathbb{Z}/p\mathbb{Z}$ with $\phi(x)=\phi(-x)$ and $\phi(\alpha x)=\alpha\phi(x)$ for all $x\in\mathbb{Z}/p\mathbb{Z}$. Let $a,b\in\mathbb{Z}/p\mathbb{Z}$ with $a\neq b$ and set $a^{\prime}:=2b-a$. Since $p>2$ we have $a\neq a^{\prime}$, but $(a,x)\cdotp(b,y)=(a^{\prime},x)\cdotp(b,y)$, a contradiction. ∎
@@
If $p=2$, one can show by computer calculation that there are no latin cycle sets of size $8$. If $p>2$ and $X$ is not simple, then we obtain a latin cycle set having size in $\{p,p^{2}\}$, but this contradicts [24, Theorem 2.13] and 56. ∎
@@
For some classes of cycle sets, the answer to the previous question is positive (see [10, 28]). In [29], Kanrar and Rump found a family of counterexamples: they are non-simple cycle sets of $2-$power size and the smaller one has size $256$. Below, we show that the construction of smaller counterexamples that are also simple cycle sets is possible. At first, we need a lemma implicitly contained in [38, Section 5].
@@
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ provided by 54 with $\lambda=1$, $\mu\neq 1$, and $c=(0,\ldots,0).$ Since the determinant of $I-B$ is $1-\mu$, by Lemma 59 every left multiplication of $X$ has only one fixed point. Let $x$ be an element of $X$. If every non-trivial cycle of $\sigma_{x}$ is not coprime with $p$, and hence divisible by $p$, we obtain that number of the fixed points of $\sigma_{x}$ is divisible by $p$, a contradiction. Therefore, $X$ is a counterexample to 1.
@@
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ obtained by 54 with $\lambda=\mu=1$ and $c:=(0,\ldots,0)$. Since $\psi$ has $-(x-1)^{p}$ as characteristic polynomial, the Jordan canonical form of $\psi$ exists in $\mathbb{Z}/p\mathbb{Z}$ and is an upper triangular matrix. Let :=$\mathcal{B}:=\{e_{1},\ldots,e_{p}\}$ the basis of $X$ that realizes the required form. Then, the subspace $H$ generated by the set $\{e_{1},\ldots,e_{p-1}\}$ is $\psi$-invariant and has dimension $p-1$, moreover $X=H\oplus<e_{p}>$ and it correspond bijectively to the set $Y\times Z$ given by $Y:=\{0\}\oplus<e_{p}>$ and $Z:=H\oplus\{0\}$, where the correspondence is given by $h+te_{p}\mapsto((0,te_{p}),(h,0))$ for all $h\in H$, $t\in\mathbb{Z}/p\mathbb{Z}$. By 34, $X$ has a complete blocks system $\{X_{y}\}_{y\in Y}$, where $X_{y}:=\{(y,z)\mid z\in Z\}$, for $y\in Y$, has size $p^{p-1}$, and with $|Y|=p$.

### v2

@@ 5.3 Simple cycle sets with regular displacement group
From now on, in this subsection we indicate by $(p,G,\phi,\psi)$ a $4$-uple such that:
@@
    $(\phi,\psi)\in Sym(G)\times Aut(G,+)$ such that Eq. 8 holds for all $x,y\in G$.
By 34, 41 and 43, we are able to construct all the simple cycle sets with nilpotent regular permutation group. We leave the proof to the reader since it follows by the cited results.
Theorem 44.
Let $(p,G,\phi,\psi)$ be such that every normal subgroup $H$ of $G$ invariant under $\psi$ is not invariant under $\phi$. Then, the pair $(X,\bullet)$ given by $X:=G$ and
$x\bullet y:=\phi(x)+\psi(y)$
for all $x,y\in G$ is a non-trivial simple cycle set with $\mathrm{Dis}(X)\cong G$.
Conversely, every simple cycle set $X$ having nilpotent regular displacement group can be constructed in this way.
If $X$ is a simple cycle set with nilpotent regular displacement group, by 22 it can be constructed by a suitable cycle set having a $p$-group permutation group.
Corollary 45.
Let $Y$ be a simple cycle set having nilpotent regular displacement group. Then, it is a deformation of a right linear cycle set $X$ by a suitable $\alpha\in Aut(X)$ such that $\mathcal{G}(X)$ is a $p$-group for some prime number $p$.
@@
It follows by 41, 43, 37, 22, and 29. ∎
By the previous corollary, if $Y$ is a simple cycle set with nilpotent regular displacement group, then it is a deformation of some cycle set $X$ by an automorphism $\alpha$. Unfortunately, we do not know if such an $X$ is simple. Moreover, we do not know if the deformation of a simple cycle set is again simple, but we are able to show this under some additional hypothesis.
Proposition 46.
Let $X$ be a simple cycle set having a $p$-group permutation group $\mathcal{G}(X)$. Let $\alpha$ be an automorphism of $X$ having order coprime with $p$ and fixing an element $x\in X$. Then, the deformation $X_{\alpha}$ of $X$ is simple.
Proof.
Suppose that $X_{\alpha}$ is not a simple cycle set. Then, by [22, Lemma 4.5] the left brace $\mathcal{G}(X_{\alpha})$ is isomorphic to $\mathcal{G}(X)\rtimes<\alpha>$, and by 39 there exist a nonzero ideal $J$ of $\mathcal{G}(X_{\alpha})$ that acts not transitively on $X_{\alpha}$. Then, the ideal $J\cap\mathrm{Dis}(X_{\alpha})$, which is equal to $J\cap\mathcal{G}(X_{\alpha})^{2}$, contains $J*\mathcal{G}(X)$, therefore if $J\cap\mathrm{Dis}(X_{\alpha})=0$ we have $J*\mathcal{G}(X_{\alpha})=0$ and hence $J\subseteq Soc(\mathcal{G}(X_{\alpha}))=0$, but this is not possible because $J$ is nonzero. Hence $J\cap\mathrm{Dis}(X_{\alpha})$ is a nonzero ideal of $\mathcal{G}(X_{\alpha})$ contained in $\mathrm{Dis}(X_{\alpha})$. Since by 29 $\mathrm{Dis}(X_{\alpha})=\mathrm{Dis}(X)$, we have that $J\cap\mathrm{Dis}(X)$ is a nonzero ideal of $\mathcal{G}(X)$ that acts not transitively on $X$, hence by 39 $X$ is not simple, a contradiction. ∎
Corollary 47.
Let $(p,G,\phi,\psi)$ be such that $\psi$ has order $p^{s}$ for some natural number $s$, and every normal subgroup $H$ of $G$ invariant under $\psi$ is not invariant under $\phi$. Moreover, let $\alpha\in Aut(G,\cdotp)$, where $x\cdotp y:=\phi(x)+\psi(y)$ for all $x,y\in G$, such that $\alpha$ has coprime order with $p$ and $\alpha(z)=z$ for some $z\in G$. Then, the pair $(X,\bullet)$ given by $X:=G$ and
$x\bullet y:=\alpha(\phi(x)+\psi(y))$
for all $x,y\in G$ is a simple cycle set.
Proof.
It follows by 37, 44 and 46. ∎
@@ 6 Affine cycle sets
The following result, that is a reformulation of [9, Theorem $4.18$], allows to detect the class of affine cycle sets by group-theoretic properties.
@@
    Let $(X,\cdotp)=\mathrm{Aff}(A,\phi,\psi,c)$ be a latin affine cycle set with underlying abelian group $(A,+)$. As showed in [9, Section $4$], for affine cycle sets Eq. 6 is equivalent to the relation $[\psi,\phi^{-1}]=id_{X}$, that is also the defining relation of the first Weyl Algebra $A_{1}(\mathbb{Z})$. Thus $(X,+)$ has a structure of $A_{1}(\mathbb{Z})$-module defined by $\rho:A_{1}(\mathbb{Z})\longrightarrow\operatorname{End}{(A,+)}$ with $\rho(a)=\psi$ and $\rho(b)=\phi^{-1}$.
@@
Note that, if $(M,+)$ is an elementary abelian $p$-group, then $(M,+)$ is a $\mathbb{Z}/p\mathbb{Z}$-module and so $\mathrm{Aff}(M,\phi,\psi,c)$ is a $A_{1}(\mathbb{Z}/p\mathbb{Z})$-module. Hence we have a representation of the algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ over $M$, and so the dimension of the representation must be a multiple of $p$ [9, Proposition 4.7].
@@ 6.1 Simple affine cycle sets
Let $p$ be a prime number. As a main consequence of 3 and 49, we obtain that every finite simple affine cycle set corresponds to a finite-dimensional irreducible representations $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ together with a constant $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, where $n:=dim(\rho)$.
@@
Conversely, suppose that $X$ is an affine simple latin cycle set. Then, by [9, Section 4] and 49 there exist a natural number $n$, a prime number $p$ dividing $n$, an element $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, and two automorphism $\phi$ and $\psi$ of $(\mathbb{Z}/p\mathbb{Z})^{n}$, such that $X:=(\mathbb{Z}/p\mathbb{Z})^{n}$ and $x\cdotp y:=\phi(x)+\psi(y)+c$ for all $x,y\in(\mathbb{Z}/p\mathbb{Z})^{n}$. Then, the assignments $a\mapsto\psi$ and $b\mapsto\phi^{-1}$ gives rise to a representation $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ (see also Remark 3). By simplicity of $X$, the irreducibility of $\rho$ follows. ∎
Remark 4.
We highlight that by 48, the simple cycle sets $X$ with displacement group abelian and normal in the total permutation group $\mathcal{TG}(X)$ are all and only the ones provided in the previous theorem.
Now, we indicate by $(p,\rho,c,\eta,g)$ a quintuple such that:
@@
    $\rho$ is a representation of the Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ with $\rho(a):=\psi$ and $\rho(b):=\phi$ invertible matrices;
@@
    $(\eta,g)\in C_{Aut(G,+)}(A,B)\times(\mathbb{Z}/p\mathbb{Z})^{n}$ is such that $\eta(c)-c=Ag+Bg-g$.
Below, we show an affine-version of 47. The proof is left to the reader, since it follows by the previous results and is standard.
Corollary 51.
Let $(p,\rho,c,\eta,g)$ be a quintuple such that $\rho$ is an irreducible representation, $\psi$ as order $p^{s}$ for some natural number $s$, and $\eta$ has order coprime with $p$. Then, the pair $((\mathbb{Z}/p\mathbb{Z})^{n},\cdotp)$ with
$x\cdotp y:=g+\eta(\rho(b)^{-1}x+\rho(a)y+c)$
@@ 6.2 Simple affine cycle sets of minimal size
By 50, several examples of simple cycle sets can be obtained by irreducible representations of the first Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$. Even if these representations are not classified over arbitrary fields (to our knowledge), several examples can be recovered from the irreducible representations of the first Weyl Algebras on algebrically closed fields, that are completely classified. Indeed, if $K$ is an algebrically closed field of characteristic $p>0$, all the finite dimensional irreducible representations of $A_{1}(K):=K[a,b]/(ab-ba-1)$ are given by the $p\times p$ matrices
@@
By 50 and 53, we are able to construct concretely all the affine simple cycle sets of size $p^{p}$, for an arbitrary prime number $p$. The next goal is to distinguish the isomorphism classes.
@@
According to 50 and 53, we have that $X$ is given by $X(\lambda,\mu,c)=\mathrm{Aff}(\mathbb{Z}_{p}^{p},(\lambda I+\Delta)^{-1},\mathcal{A}_{\mu},c)$. If $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic then the matrices $A_{\mu}$ and $A_{\mu^{\prime}}$ (resp. $\mu I+\Delta$ and $\mu^{\prime}I+\Delta$) are conjugate according to 56. Therefore such matrices have the same eigenvalues, i.e. $\lambda=\lambda^{\prime}$ (resp. $\mu=\mu^{\prime}$). By this fact and 56, $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic if and only if and only if $\lambda=\lambda^{\prime}$, $\mu=\mu^{\prime}$ and there exists $\alpha\in GL_{p}(\mathbb{Z}/p\mathbb{Z})$ and $u\in\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$ such that $\mu I+\Delta=\alpha(\mu I+\Delta)\alpha^{-1}$, $\mathcal{A}_{\mu}=\alpha\mathcal{A}_{\mu}\alpha^{-1}$ and $c^{\prime}=\alpha(c+u)$. By virtue of Lemma 55 then $\alpha=sI$ with $s\neq 0$. Therefore
@@ 7 Applications and examples
We start by a non-existence result that extend [9, Lemma 4.13]. In particular, we show that we have no latin cycle sets with cyclic displacement group.
@@
Let $n:=|X|$ and suppose that $\mathrm{Dis}(X)$ is a cyclic group. At first, suppose that $n$ has an odd prime divisor $p$ and let $m$ be the positive integer such that $n/m=p$. As in 42, the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$ and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $H$ induces a congruence $\sim$ on $X$ such that $|X/\sim|=p$, and this implies that there exist a latin cycle set of size $p$, but this contradicts [25, Theorem 2.13]. Now, we consider the case $n=2^{t}$ for some natural number $t$. The case $t=1$ is not possible by [25, Theorem 2.13]. The case $t=2$ is excluded by an inspection of small cycle sets. The case $t>2$ can be showed as in the first part by [2, Proposition 5.4]. ∎
By [22] every indecomposable and irretractable cycle set of size $p^{2}$ is simple. A complete description of the latin ones follows by [22, Theorem $5.1$], as we show in the following result.
@@
By 3, $X$ must be indecomposable and irretractable. If $p=2$, we know by [9] that the two indecomposable and irretractable cycle sets of order $4$ are latin. If $p>2$, by [22, Theorem $5.1$] $X$ is given by $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ and $(a,x)\cdotp(b,y)=(\alpha b+\alpha x,\alpha y+\alpha\phi(b-a))$, where $\alpha\in\mathbb{Z}/p\mathbb{Z}\setminus\{0\}$ and $\phi$ is a map from $\mathbb{Z}/p\mathbb{Z}$ to $\mathbb{Z}/p\mathbb{Z}$ with $\phi(x)=\phi(-x)$ and $\phi(\alpha x)=\alpha\phi(x)$ for all $x\in\mathbb{Z}/p\mathbb{Z}$. Let $a,b\in\mathbb{Z}/p\mathbb{Z}$ with $a\neq b$ and set $a^{\prime}:=2b-a$. Since $p>2$ we have $a\neq a^{\prime}$, but $(a,x)\cdotp(b,y)=(a^{\prime},x)\cdotp(b,y)$, a contradiction. ∎
@@
If $p=2$, one can show by computer calculation that there are no latin cycle sets of size $8$. If $p>2$ and $X$ is not simple, then we obtain a latin cycle set having size in $\{p,p^{2}\}$, but this contradicts [25, Theorem 2.13] and 59. ∎
@@
For some classes of cycle sets, the answer to the previous question is positive (see [11, 29]). In [30], Kanrar and Rump found a family of counterexamples: they are non-simple cycle sets of $2-$power size and the smaller one has size $256$. Below, we show that the construction of smaller counterexamples that are also simple cycle sets is possible. At first, we need a lemma implicitly contained in [39, Section 5].
@@
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ provided by 57 with $\lambda=1$, $\mu\neq 1$, and $c=(0,\ldots,0).$ Since the determinant of $I-B$ is $1-\mu$, by Lemma 62 every left multiplication of $X$ has only one fixed point. Let $x$ be an element of $X$. If every non-trivial cycle of $\sigma_{x}$ is not coprime with $p$, and hence divisible by $p$, we obtain that number of the fixed points of $\sigma_{x}$ is divisible by $p$, a contradiction. Therefore, $X$ is a counterexample to 1.
@@
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ obtained by 57 with $\lambda=\mu=1$ and $c:=(0,\ldots,0)$. Since $\psi$ has $-(x-1)^{p}$ as characteristic polynomial, the Jordan canonical form of $\psi$ exists in $\mathbb{Z}/p\mathbb{Z}$ and is an upper triangular matrix. Let :=$\mathcal{B}:=\{e_{1},\ldots,e_{p}\}$ the basis of $X$ that realizes the required form. Then, the subspace $H$ generated by the set $\{e_{1},\ldots,e_{p-1}\}$ is $\psi$-invariant and has dimension $p-1$, moreover $X=H\oplus<e_{p}>$ and it correspond bijectively to the set $Y\times Z$ given by $Y:=\{0\}\oplus<e_{p}>$ and $Z:=H\oplus\{0\}$, where the correspondence is given by $h+te_{p}\mapsto((0,te_{p}),(h,0))$ for all $h\in H$, $t\in\mathbb{Z}/p\mathbb{Z}$. By 34, $X$ has a complete blocks system $\{X_{y}\}_{y\in Y}$, where $X_{y}:=\{(y,z)\mid z\in Z\}$, for $y\in Y$, has size $p^{p-1}$, and with $|Y|=p$.